# thanks: Corresponding authorthanks: Corresponding author

### v1

@@ Abstract
Recently, the baryon acoustic oscillations (BAO) measurements from the DESI survey have suggested hints of dynamical dark energy, challenging the standard $\Lambda$CDM model. In this work, we adopt an improved inverse distance ladder approach based on the latest cosmological data to provide a model-independent perspective, employing a global parametrization based on cosmic age (PAge). Our analysis incorporates DESI DR2 BAO measurements, cosmic chronometer (CC) data, and type Ia supernovae (SNe) observations from either the DESY5 or PantheonPlus datasets. For the DESY5+DESI DR2+CC datasets, we obtain $H_{0}=67.91\pm 2.33~{}\mathrm{km~{}s^{-1}~{}Mpc^{-1}}$. This value is consistent with the Planck 2018 result, while shows $2.0\sigma$ tension with the SH0ES measurement. Furthermore, by mapping specific cosmological models into PAge approximation parameter space $(p_{\mathrm{age}},\eta)$, our model-independent analysis reveals a notable deviation from the $\Lambda\mathrm{CDM}$ model, as indicated by the DESY5 and DESI DR2 datasets. Finally, DESY5+DESI DR2+CC datasets provide nearly decisive evidence favoring the PAge model over the standard $\Lambda\mathrm{CDM}$ model. These findings highlight the need for further investigation into the expansion history to better understand the deviations from the $\Lambda\mathrm{CDM}$ model.
@@ I INTRODUCTION
The simplest $\Lambda$ cold dark matter ($\Lambda\mathrm{CDM}$) model has been remarkably consistent with most astronomical observations across different epochs and scales of the universe Riess et al. (1998); Perlmutter et al. (1999); Brout et al. (2022); Asgari et al. (2021); Troxel et al. (2018); Aiola et al. (2020). However, the persistent discrepancy in the value of the Hubble constant $H_{0}$ inferred from the early and late universe, known as the Hubble tension, is challenging the $\Lambda\mathrm{CDM}$ model Di Valentino et al. (2021a); Perivolaropoulos and Skara (2022); Shah et al. (2021); Verde et al. (2019). Among various discrepancies, the most significant is the over $5\sigma$ divergence between the Planck 2018 estimate of $H_{0}=67.4\pm 0.5\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, indirectly inferred from the cosmic microwave background (CMB) data assuming the $\Lambda\mathrm{CDM}$ model Aghanim et al. (2020), and the SH0ES result of $H_{0}=73.04\pm 1.04\,\mathrm{km\ s^{-1}\ Mpc^{-1}}$ directly measured from the type Ia supernovae (SNe) calibrated using the Cepheid variables Riess et al. (2022). The tension cannot be merely explained by a statistical fluctuation, even after the reanalysis of systematic uncertainties Akrami et al. (2020); Zhang et al. (2017a). Therefore, the irreconcilable tension may indicate underlying systematics or new physics beyond $\Lambda\mathrm{CDM}$ Di Valentino et al. (2021b); Bernal et al. (2016); Abdalla et al. (2022).
@@
In early-time solutions, new physics is introduced to reduce the sound horizon, $r_{\mathrm{d}}$, thereby increasing $H_{0}$ while preserving the baryon acoustic oscillation (BAO) measurement of $r_{\mathrm{d}}H_{0}$. One realization involves modifying expansion history by introducing additional energy components around the epoch of matter-radiation equality, such as early dark energy Poulin et al. (2019); Karwal and Kamionkowski (2016); Mörtsell and Dhawan (2018); Kamionkowski et al. (2014); Berghaus and Karwal (2020) or extra neutrino species Zhang et al. (2015); Zhao et al. (2017a, b); Feng et al. (2020); Kreisch et al. (2020); Wyman et al. (2014); Sakstein and Trodden (2020). Alternatively, early-time solutions may also involve modifying the recombination history Sekiguchi and Takahashi (2021); Chiang and Slosar (2018); Mirpoorian et al. (2024) or introducing primordial magnetic fields Jedamzik and Pogosian (2020); Jedamzik et al. (2025); Jedamzik and Abel (2013). However, these early-time solutions are in tension with other cosmological observables, such as the fluctuation amplitude $\sigma_{8}$ and the integrated Sachs-Wolfe effect, as discussed in the no-go arguments of Refs. Hill et al. (2020); Krishnan et al. (2020); Jedamzik et al. (2021); Vagnozzi (2023); Dutta et al. (2019). Late-time solutions modify the Hubble parameter $H(z)$ in the late universe, so that it can accommodate the SH0ES $H_{0}$ result locally while leaving the CMB anisotropy spectrum unchanged, such as the deformation of the dark-energy equation of state Wang et al. (2016); Agrawal et al. (2023); Li and Shafieloo (2019); Li (2004); Shapiro and Sola (2002), modified gravity Aluri et al. (2023); Capozziello and De Laurentis (2011); De Felice and Tsujikawa (2010), parameter transitions in the late universe Mortonson et al. (2009); Kazantzidis and Perivolaropoulos (2019). Analogous to early-time solutions, various no-go arguments have also been raised against the late-time scenarios. The requirement of consistency between low redshift cosmological distance measurements from BAO and SNe imposes strong constraints on the late-time solutions Benevento et al. (2020); Yang et al. (2021). Moreover, these solutions cannot reconcile the $H_{0}$ tension and $\sigma_{8}$ tension simultaneously, as Refs. Guo et al. (2019); Gao et al. (2021); Heisenberg et al. (2022); Dutta et al. (2019); Alestas and Perivolaropoulos (2021) pointed out.
@@
In light of the plethora of proposed cosmological models and corresponding no-go arguments, it is compelling to explore new and independent measurements of $H_{0}$. For instance, techniques such as strong gravitational lensing time delay Birrer et al. (2019); Shajib et al. (2020) and gravitational wave standard sirens Guidorzi et al. (2017); Song et al. (2025) provide an independent perspective on this tension. We can also employ other independent calibrators to replace Cepheid variables in the distance ladder, such as megamasers Pesce et al. (2020), the tip of the red giant branch Freedman (2021), surface brightness fluctuations Blakeslee et al. (2021), Mira variables Huang et al. (2019).
@@
The inverse distance ladder (IDL) calibrates the SNe via anchored BAO instead of Cepheid variables, providing a model-insensitive method to infer $H_{0}$ Aubourg et al. (2015); Cuesta et al. (2015); Verde et al. (2017); Lemos et al. (2019). In the traditional IDL, the BAO data are anchored using an $r_{\mathrm{d}}$ prior derived from the CMB or big bang nucleosynthesis. Then the data are used to calibrate SNe to give an absolute distance. Finally, the distance is extrapolated to $z=0$ to infer $H_{0}$ under a fiducial $H(z)$ model Percival et al. (2010); Heavens et al. (2014); Abbott et al. (2018); Alam et al. (2021); Camilleri et al. (2025); Adame et al. (2025); Macaulay et al. (2019).
@@
It should be noted that the $r_{\mathrm{d}}$ prior introduces dependence on early-universe physics Aylor et al. (2019). The effect of this $r_{\mathrm{d}}$ prior has been systematically evaluated in Refs. Barua and Desai (2024); Luongo and Muccino (2024). To mitigate such dependence, the $r_{\mathrm{d}}$ prior can be replaced with other independent measurements. Cosmic chronometer (CC) measure $H(z)$ directly, rather than through integrals of $H(z)$, such as those involved in lensing distances from strong lensing systems Wojtak and Agnello (2019); Taubenberger et al. (2019); Wang et al. (2022); Gong et al. (2024), thus offering higher sensitivity to cosmological parameters Moresco (2024); Vagnozzi et al. (2022); Jimenez et al. (2019). CC data have been widely applied as an effective replacement for $r_{\mathrm{d}}$ in constructing IDL Guo et al. (2025); Luongo and Muccino (2024); Favale et al. (2023); Cai et al. (2022a).
@@
Furthermore, adopting a fiducial $H(z)$ model will inevitably introduce model dependence. To avoid this, the traditional approach is to parameterize $H(z)$ via a Taylor expansion in terms of $z$ or $y\equiv 1-a=z/(1+z)$, known as the cosmography method Camilleri et al. (2025); Pourojaghi et al. (2025); Luongo and Muccino (2024); Lemos et al. (2019); Cattoen and Visser (2007); Zhang et al. (2017b); Chiba and Nakamura (1998); Capozziello et al. (2011); Aviles et al. (2012). Alternatively, purely data-driven techniques such as Gaussian Process (GP) have been employed to reconstruct $H(z)$ Guo et al. (2025); Jiang et al. (2024); He et al. (2022); Mukherjee and Sen (2024); Yang et al. (2023). However, both approaches possess intrinsic limitations. The Taylor expansion is problematic at $z\sim 1$ Banerjee et al. (2021); Cai et al. (2022a, b); Zhang et al. (2017b). GP methods are sensitive to the choice of kernel functions and tend to yield smaller errors at the cost of strongly correlated results Ó Colgáin and Sheikh-Jabbari (2021); Kjerrgren and Mortsell (2022). To address these limitations and achieve a fully model-independent method, we adopt a global parametrization based on the cosmic age (PAge) model Huang (2020); Luo et al. (2020) to accurately reproduce late-time cosmological behavior across a wide redshift range.
@@
The PAge model can be interpreted as an effective description of homogeneous late-time scenarios. Moreover, for models where the $H(z)$ evolves rapidly, a Taylor expansion in terms of $z$ or $y$ requires high-order terms to accurately capture the behavior. However, since the cosmic age $t$ is defined through the integral of $H(z)$, it evolves more smoothly. As a result, even pronounced changes in $H(z)$ can typically be captured by a low-order expansion in $t$. So the PAge model has been employed as a substitute for specific physical models to mitigate model dependence. The studies Huang et al. (2021a, 2022); Li et al. (2022) have employed the PAge model to analyze gamma-ray bursts, redshift-space distortions and quasars. Refs. Cai et al. (2022a, b); Huang et al. (2025a) have derived no-go theorems against late-time solutions through the comparison between the PAge (denoted as late-time scenarios) and $\Lambda\mathrm{CDM}$ models.
@@
Recently, the Dark Energy Spectroscopic Instrument (DESI) collaboration reported a $3.9\sigma$ hint of dynamical dark energy in the first data release (DR1), which was further increased to $4.2\sigma$ in the second data release (DR2) Adame et al. (2025); Abdul Karim et al. (2025). This finding has sparked intense discussions about possible new physics and systematic uncertainties Pourojaghi et al. (2024); Colgáin and Sheikh-Jabbari (2024); Jia et al. (2025); Wang et al. (2024a); Huang et al. (2025b); Cortês and Liddle (2024); Efstathiou (2024); Abreu and Turner (2025); Pang et al. (2024); Fikri et al. (2024); Feng et al. (2025); Du et al. (2025a); Li et al. (2024a, b, 2025); Du et al. (2025b); Pang et al. (2025); Wang and Piao (2025, 2024); Lodha et al. (2025); Pan et al. (2025); You et al. (2025); Pan and Ye (2025); Silva et al. (2025). This behavior, which deviates from $\Lambda\mathrm{CDM}$ can be effectively described by the PAge model. This motivates us to propose a PAge-improved IDL to yield fully model-independent cosmological constraints in this work. Our results will provide model-independent perspectives on the $H_{0}$ tension and the hint of dynamical dark energy.
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the PAge model and cosmological datasets used in our analysis. Section III presents the cosmological constraints and related discussion. The paper concludes with a summary in Section IV.
@@ II METHODOLOGY
II.1 The PAge model
In IDL, a fiducial model is needed to trace the universal expansion history $H(z)$ back to $z=0$ to determine $H_{0}$. To achieve a model-independent analysis of late-time data, the simplest technique is a Taylor expansion Visser (2004); Cattoen and Visser (2007) about redshift $z$ Zhang et al. (2017b); Chiba and Nakamura (1998) or y-redshift, $y\equiv 1-a=z/(z+1)$ Capozziello et al. (2011); Aviles et al. (2012). However, both ways become problematic at $z\sim 1$ due to the deviation from the target model, as discussed in Refs. Hu and Wang (2022); Cai et al. (2022a).
The PAge model is introduced to provide a global approximation of $H(z)$ with the cosmic age $t$ Huang (2020). It remains consistent with the target model at high redshift and is more precise than $z$- or $y$-expansion. Actually, it is a calibrated Taylor expansion around $t$. Motivated by the fact that the product of the Hubble constant and current age of the universe $H_{0}t_{0}\sim\mathcal{O}(1)$ differs little among most cosmological models Vagnozzi et al. (2022), we begin by expanding $H(t)t$ about the current cosmic age $t_{0}$ as Wang et al. (2024b)
$H(t)t=H_{0}t_{0}+H_{0}(t-t_{0})+\left.\dot{H}\,\right|_{t=t_{0}}(t-t_{0})t+\ldots.$ (1)
Neglecting the brief radiation-dominated era, matter domination in the early universe requires Eq. (1) to satisfy the following condition
$\lim\limits_{t\to 0}H(t)t=\frac{2}{3},$ (2)
which introduces a calibration to Taylor expansion to prevent it from diverging at high redshift. Moreover, for simplicity and Occam’s razor, we only expand $H(t)t$ to the quadratic order. The higher-order terms have been proven to be unnecessary given the accuracy level of current data in Refs. Cai et al. (2022b); Huang et al. (2021b). Thus, the final calibrated expansion takes the following form

### v2

@@ Abstract
[…]
@@ I INTRODUCTION
The simplest $\Lambda$ cold dark matter ($\Lambda\mathrm{CDM}$) model has been remarkably consistent with most astronomical observations across different epochs and scales of the Universe Riess et al. (1998); Perlmutter et al. (1999); Brout et al. (2022); Asgari et al. (2021); Troxel et al. (2018); Aiola et al. (2020). However, the persistent discrepancy in the value of the Hubble constant $H_{0}$ inferred from the early and late Universe, known as the Hubble tension, is challenging the $\Lambda\mathrm{CDM}$ model Di Valentino et al. (2021a); Perivolaropoulos and Skara (2022); Shah et al. (2021); Verde et al. (2019). Among various discrepancies, the most significant is the over $5\sigma$ divergence between the Planck 2018 estimate of $H_{0}=67.4\pm 0.5\,\mathrm{km\,s^{-1}\,Mpc^{-1}}$, indirectly inferred from the cosmic microwave background (CMB) data assuming the $\Lambda\mathrm{CDM}$ model Aghanim et al. (2020), and the SH0ES result of $H_{0}=73.04\pm 1.04\,\mathrm{km\ s^{-1}\ Mpc^{-1}}$ directly measured from the type Ia supernovae (SNe) calibrated using the Cepheid variables Riess et al. (2022). The tension cannot be merely explained by a statistical fluctuation, even after the reanalysis of systematic uncertainties Akrami et al. (2020); Zhang et al. (2017a). Therefore, the irreconcilable tension may indicate underlying systematics or new physics beyond $\Lambda\mathrm{CDM}$ Di Valentino et al. (2021b); Bernal et al. (2016); Abdalla et al. (2022); Di Valentino et al. (2025).
@@
In early-time solutions, new physics is introduced to reduce the sound horizon, $r_{\mathrm{d}}$, thereby increasing $H_{0}$ while preserving the baryon acoustic oscillation (BAO) measurement of $r_{\mathrm{d}}H_{0}$. One realization involves modifying expansion history by introducing additional energy components around the epoch of matter-radiation equality, such as early dark energy Poulin et al. (2019); Karwal and Kamionkowski (2016); Mörtsell and Dhawan (2018); Kamionkowski et al. (2014); Berghaus and Karwal (2020) or extra neutrino species Zhang et al. (2015a); Zhao et al. (2017a, b); Feng et al. (2020); Kreisch et al. (2020); Wyman et al. (2014); Sakstein and Trodden (2020); Zhang et al. (2014a). Alternatively, early-time solutions may also involve modifying the recombination history Sekiguchi and Takahashi (2021); Chiang and Slosar (2018); Mirpoorian et al. (2025) or introducing primordial magnetic fields Jedamzik and Pogosian (2020); Jedamzik et al. (2025); Jedamzik and Abel (2013). However, these early-time solutions are in tension with other cosmological observables, such as the fluctuation amplitude $\sigma_{8}$ [^1] However, using $\sigma_{8}$ may bias the comparison between models with high $H_{0}$ and $\Lambda\mathrm{CDM}$, since the reference scale $R=8h^{-1}~\mathrm{Mpc}$ depends on $h$, which changes the interpretation of $\sigma_{8}$ Forconi et al. (2025); Sanchez (2020). and the integrated Sachs-Wolfe effect, as discussed in the no-go arguments of Refs. Hill et al. (2020); Krishnan et al. (2020); Jedamzik et al. (2021); Vagnozzi (2023); Dutta et al. (2019). Late-time solutions modify the Hubble parameter $H(z)$ in the late universe, so that it can accommodate the SH0ES $H_{0}$ result locally while leaving the CMB anisotropy spectrum unchanged, such as the deformation of the dark-energy equation of state Wang et al. (2016); Agrawal et al. (2023); Li and Shafieloo (2019); Li (2004); Shapiro and Sola (2002); Zhang et al. (2012); Li et al. (2013); Zhang et al. (2015b); Feng and Zhang (2016), modified gravity Aluri et al. (2023); Capozziello and De Laurentis (2011); De Felice and Tsujikawa (2010), and parameter transitions in the late Universe Mortonson et al. (2009); Kazantzidis and Perivolaropoulos (2019). Analogous to early-time solutions, various no-go arguments have also been raised against the late-time scenarios. The requirement of consistency between low redshift cosmological distance measurements from BAO and SNe imposes strong constraints on the late-time solutions Benevento et al. (2020); Yang et al. (2021). Moreover, these solutions cannot reconcile the $H_{0}$ tension and $S_{8}$ tension [^2] The $S_{8}$ tension is alleviated as indicated by the updated Kilo-Degree Survey cosmic shear constraints with improved redshift distribution estimation and calibration, which found consistency ($0.73\sigma$) with Planck Wright et al. (2025). simultaneously, as Refs. Guo et al. (2019); Gao et al. (2021); Heisenberg et al. (2022); Dutta et al. (2019); Alestas and Perivolaropoulos (2021) pointed out.
@@
In light of the plethora of proposed cosmological models and corresponding no-go arguments, it is compelling to explore new and independent measurements of $H_{0}$. For instance, techniques such as strong gravitational lensing time delay Birrer et al. (2019); Shajib et al. (2020) and gravitational wave standard sirens Guidorzi et al. (2017); Song et al. (2025); Jin et al. (2025) provide an independent perspective on this tension. We can also employ other independent calibrators to replace Cepheid variables in the distance ladder, such as megamasers Pesce et al. (2020), the tip of the red giant branch Freedman (2021), surface brightness fluctuations Blakeslee et al. (2021), and Mira variables Huang et al. (2019).
@@
The inverse distance ladder (IDL) calibrates the SNe via anchored BAO instead of Cepheid variables, providing a model-insensitive method to infer $H_{0}$ Aubourg et al. (2015); Cuesta et al. (2015); Verde et al. (2017); Lemos et al. (2019); Yu et al. (2018); Gómez-Valent and Amendola (2018); Haridasu et al. (2018); Gómez-Valent et al. (2024). In the traditional IDL, the BAO data are anchored using an $r_{\mathrm{d}}$ prior derived from the CMB or big bang nucleosynthesis. Then the data are used to calibrate SNe to give an absolute distance. Finally, the distance is extrapolated to $z=0$ to infer $H_{0}$ under a fiducial $H(z)$ model Percival et al. (2010); Heavens et al. (2014); Abbott et al. (2018); Alam et al. (2021); Camilleri et al. (2025); Adame et al. (2025); Macaulay et al. (2019).
@@
It should be noted that the $r_{\mathrm{d}}$ prior introduces dependence on early-Universe physics Aylor et al. (2019). The effect of this $r_{\mathrm{d}}$ prior has been systematically evaluated in Refs. Barua and Desai (2025); Luongo and Muccino (2024). To mitigate such dependence, the $r_{\mathrm{d}}$ prior can be replaced with other independent measurements. The cosmic chronometer (CC) measures $H(z)$ directly, rather than through integrals of $H(z)$, such as those involved in lensing distances from strong lensing systems Wojtak and Agnello (2019); Taubenberger et al. (2019); Wang et al. (2022); Gong et al. (2024), thus offering higher sensitivity to cosmological parameters Moresco (2024); Vagnozzi et al. (2022); Jimenez et al. (2019). CC data have been widely applied as an effective replacement for $r_{\mathrm{d}}$ in constructing IDL Guo et al. (2025); Luongo and Muccino (2024); Favale et al. (2023); Cai et al. (2022a).
@@
Furthermore, adopting a fiducial $H(z)$ derived from a specific cosmological model in IDL will inevitably introduce model dependence. To avoid this, the traditional approach is to parametrize $H(z)$ via a Taylor expansion in terms of $z$ or $y\equiv 1-a=z/(1+z)$, known as the cosmography method Camilleri et al. (2025); Pourojaghi et al. (2025); Luongo and Muccino (2024); Lemos et al. (2019); Cattoen and Visser (2007); Zhang et al. (2017b); Chiba and Nakamura (1998); Capozziello et al. (2011); Aviles et al. (2012). Alternatively, purely data-driven techniques such as the Gaussian process (GP) have been employed to reconstruct $H(z)$ Guo et al. (2025); Jiang et al. (2024); He et al. (2022); Mukherjee and Sen (2024); Yang et al. (2023). However, both approaches possess intrinsic limitations. The Taylor expansion is problematic at $z\sim 1$ Banerjee et al. (2021); Cai et al. (2022a, b); Zhang et al. (2017b). GP methods are sensitive to the choice of kernel functions and tend to yield smaller errors at the cost of strongly correlated results Ó Colgáin and Sheikh-Jabbari (2021); Kjerrgren and Mortsell (2022). To address these limitations, we adopt a global parametrization based on the cosmic age (PAge) Huang (2020); Luo et al. (2020) and its extension, a more accurate parametrization based on cosmic age (MAPAge) Cai et al. (2022b); Huang et al. (2021a), to accurately reproduce late-time cosmological behaviors.
@@
The PAge model can be interpreted as an effective description of homogeneous late-time scenarios. The studies Huang et al. (2021b, 2022); Li et al. (2022) have employed the PAge model to analyze gamma-ray bursts, redshift-space distortions, and quasars. References Cai et al. (2022a, b); Huang et al. (2025a) have derived no-go theorems against late-time solutions through the comparison between the PAge (denoted as late-time scenarios) and $\Lambda\mathrm{CDM}$ models.
@@
Recently, the Dark Energy Spectroscopic Instrument (DESI) Collaboration reported a $3.9\sigma$ hint of dynamical dark energy in the first data release (DR1), which was further increased to $4.2\sigma$ in the second data release (DR2) Adame et al. (2025); Abdul Karim et al. (2025). This finding has sparked intense discussions about possible new physics and systematic uncertainties Malekjani et al. (2025); Colgáin and Sheikh-Jabbari (2024); Jia et al. (2025); Wang et al. (2024a); Huang et al. (2025b); Cortês and Liddle (2024); Efstathiou (2025); Abreu and Turner (2025); Pang et al. (2025a); Fikri et al. (2025); Feng et al. (2025a); Du et al. (2025a); Li et al. (2025a, 2024, b); Du et al. (2025b); Pang et al. (2025b); Wang et al. (2025a); Wang and Piao (2024); Lodha et al. (2025); Pan et al. (2025); You et al. (2025); Pan and Ye (2025); Silva et al. (2025); Qiang et al. (2025); Cai et al. (2025); Cline and Muralidharan (2025); Rodrigues et al. (2025); Li and Wang (2025); Bhattacharjee et al. (2025); Gialamas et al. (2025a); Wang and Piao (2025); Cline and Muralidharan (2025); Li et al. (2025c); Li and Zhang (2025); Li et al. (2025d, e); González-Fuentes and Gómez-Valent (2025); Gómez-Valent and Solà Peracaula (2025); Wang et al. (2025b); Du et al. (2025c); Feng et al. (2025b). This behavior, which deviates from $\Lambda\mathrm{CDM}$ can be effectively described by the PAge model. This motivates us to propose a PAge-improved IDL to yield fully model-independent cosmological constraints in this work. Our results will provide model-independent perspectives on the $H_{0}$ tension and the hint of dynamical dark energy.
@@
The remainder of this paper is organized as follows. In Sec. II, we introduce the PAge model and cosmological datasets used in our analysis. We present the cosmological constraints and related discussion in Sec. III. The paper concludes with a summary in Sec. IV.
@@ II METHODOLOGY
II.1 The PAge and MAPAge parametrization
Within the IDL, the PAge (MAPAge) parametrization is used to extrapolate $H(z)$ to $z=0$, thereby determining $H_{0}$ in a cosmological-model-independent manner. The PAge (MAPAge) rests on two assumptions:
  1. (i)
    At $z\gg 1$, the Universe is matter dominated, and the brief radiation-dominated era is neglected.
  2. (ii)
    In PAge, the product of the Hubble parameter $H$ and the cosmic time $t$, can be approximated by a quadratic function of $t$. In MAPAge, this product can be approximated by a cubic function of $t$.
These assumptions are independent of any specific cosmological model, such as dark energy models and modified gravity theories Huang (2020); Cai et al. (2022a, b); Huang et al. (2025a); Wang et al. (2024b). Rather than parametrizing the dark energy equation of state, PAge directly parametrizes the cosmic expansion history, providing a more straightforward probe of that history than equation of state based approaches such as $w_{0}w_{a}\mathrm{CDM}$. Compared with traditional parametrizations, such as expansions in $z$ or $y$ Capozziello et al. (2011); Zhang et al. (2017b); Visser (2004), PAge (MAPAge) can parametrize $H(z)$ over a wider redshift range with higher accuracy. As shown in Fig. 1, PAge (MAPAge) can approximate the target $w_{0}w_{a}\mathrm{CDM}$ model up to $z\sim 10^{3}$, whereas $z$ and $y$ expansion develop deviation at $z>1$ Hu and Wang (2022); Cai et al. (2022a). Moreover, for target models where $H(z)$ evolves rapidly, a Taylor expansion in terms of $z$ or $y$ requires high-order terms to accurately capture the behavior. However, since $t$ is defined through the integral of $H(z)$, it evolves more smoothly. As a result, even pronounced changes in $H(z)$ can typically be captured by a low-order expansion in $t$. Thus, the PAge (MAPAge) model serves as a global, model-independent parametrization that can be effectively adopted in IDL.
In PAge, we first rewrite $H(t)t$ as
$H(t)t=H(t)(t-t_{0})+H(t)t_{0},$ (1)
where $t_{0}$ denotes the current cosmic age. Then we perform a Taylor expansion of $H(t)$ around $t_{0}$,
$H(t)=H_{0}+\left.\dot{H}\,\right|_{t=t_{0}}(t-t_{0})+\ldots.$ (2)
Substituting Eq. (2) into Eq. (1) yields
$H(t)t=$ $H_{0}t_{0}+H_{0}(t-t_{0})(1-H_{0}t_{0}-H_{0}t_{0}q_{0})$
$-H^{2}_{0}(t-t_{0})^{2}(1+q_{0}-\frac{H_{0}t_{0}}{2}(2+3q_{0}+j_{0}))+\dots.$ (3)
The deceleration and jerk parameters, $q_{0}$ and $j_{0}$, arise from the Taylor expansion of $H(t)$ in Eq. (2). As required by assumption (ii), we keep terms up to quadratic order in Eq. (3). To meet the assumption (i), $H(t)t=2/3$ in the limit $t\to 0$, we introduce the following modification of our expansion,
$\lim\limits_{t\to 0}H(t)t=H_{0}t_{0}+H_{0}(0-t_{0})(1-H_{0}t_{0}-H_{0}t_{0}q_{0})-AH^{2}_{0}t^{2}_{0},$ (4)
where $A$ is an undetermined parameter. The limit $\lim\limits_{t\to 0}H(t)t=2/3$ is met only if $A=1+q_{0}-\frac{2}{3H_{0}^{2}t^{2}_{0}}$. Therefore, the final expansion becomes
$H(t)t=$ $H_{0}t_{0}+H_{0}(t-t_{0})(1-H_{0}t_{0}-H_{0}t_{0}q_{0})$
$-H^{2}_{0}(t-t_{0})^{2}(1+q_{0}-\frac{2}{3H_{0}^{2}t^{2}_{0}}).$ (5)
After setting
$p_{\mathrm{age}}=H_{0}t_{0},~\eta=1-\frac{3}{2}p^{2}_{\mathrm{age}}(1+q_{0}),$ (6)
the final expansion is written as